OpenStudy (sjg13e):

Improper Integral, (2 to infinity)

OpenStudy (sjg13e):

\[\int\limits_{2}^{\infty}\frac{ 1 }{ v^2 + 8v - 9 }dv\] Does it converge or diverge?

OpenStudy (sjg13e):

\[\int\limits_{}^{}\frac{ 1 }{ v^2 + 8v - 9 }dv = \frac{ -1 }{ v } + \frac{ 1 }{ 8 }\ln|v| - \frac{ 1 }{ 9 }v\] (this is what I got)

OpenStudy (sjg13e):

\[\lim_{t \rightarrow \infty}[\frac{ -1 }{ v }+\frac{ 1 }{ 8 }\ln|v|-\frac{ 1 }{ 9 }v]\] where a = 2 and b = t

OpenStudy (sjg13e):

I got \[\frac{ 1 }{ 3 }-\frac{ 1 }{ 8 }\ln|2|\] as my answer. But it's wrong. I know it converges, maybe I messed up with inputting the values?

OpenStudy (solomonzelman):

Suppose you want to know whether \(\large\color{black}{\displaystyle\sum_{n=7}^{\infty}\frac{1}{n^2} }\) converges or not. you know that converges, (to π²/6 by Euler) And a bigger series \(\large\color{black}{\displaystyle\sum_{n=7}^{\infty}\frac{1}{n^2+8v-9} }\) would all the more so converge

OpenStudy (solomonzelman):

So if the first series converges (and must be - so does its integral) THEN, all the more so the second series (and must be - so does the integral).

OpenStudy (solomonzelman):

That is ian observation if you studied about the *integral test*... have you?

OpenStudy (sjg13e):

Okay, yes. That is also expressed by the comparison test, right?

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

Comparison test, to a larger series.

OpenStudy (solomonzelman):

It should really start from n=2, not n=7, but that doesn't matter.

OpenStudy (solomonzelman):

If you want, we can solve the integral tho

OpenStudy (sjg13e):

okay thank you, id be awesome if you helped me solve it

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{2}^{\infty }\frac{1}{v^2+8v-9}{~\rm dv} }\) \(\large\color{black}{\displaystyle\int\limits_{2}^{\infty }\frac{1}{v^2+8v+16-16-9}{~\rm dv} }\) \(\large\color{black}{\displaystyle\int\limits_{2}^{\infty }\frac{1}{(v+4)^2-25}{~\rm dv} }\) set u=v+4 this is where I would start

OpenStudy (freckles):

I think I see what you did... you think \[\frac{1}{a+b+c}=\frac{1}{a}+\frac{1}{b}+\frac{1}{c} \text{ but this is not true }\]

OpenStudy (freckles):

you could use trig sub or do partial fractions

OpenStudy (solomonzelman):

yes, partical fractions from the beginning is better I guess, because with my way you still have partial fractions

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{2}^{\infty }\frac{1}{(v+9)(v-1)}{~\rm dv} }\)

OpenStudy (solomonzelman):

and do the partial fractions...

OpenStudy (freckles):

@SolomonZelman I think you could have done trig sub without partial fractions I was just offering another route

OpenStudy (solomonzelman):

trig after u sub?

OpenStudy (freckles):

\[\int\limits \frac{1}{(v+4)^2-25} dv \\ \int\limits \frac{1}{25[\frac{1}{25}(v+4)^2-1]} dv \\ \frac{1}{25} \int\limits \frac{1}{(\frac{v+4}{5})^2-1} dv \\ \text{ recall } \tan^2(x)=\sec^2(x)-1 \\ \text{ so use sub } \frac{v+4}{5}=\sec(x) \\ \frac{1}{5} dv=\sec(x) \tan(x) dx \\ \frac{1}{25} \int\limits \frac{1}{\sec^2(x)-1} 5 \sec(x) \tan(x) dx \\ \frac{1}{25} \int\limits \frac{5 \sec(x)\tan(x)}{\tan^2(x)}dx\] unless... I did something wrong we don't need partial fractions for the trig route here

OpenStudy (sjg13e):

Sorry, my wifi was spotty. I'll read through your explanation now

OpenStudy (freckles):

though I much prefer the partial fractions

OpenStudy (solomonzelman):

Yeah, definitely...